systms of linear equations an problem solving the substituition method
3s-4t=14
5s+t=8

Answer by elima(1433)

3s-4t=14
5s+t=8
Lets start with the 2nd equation;
5s+t=8
solve for t;
t=8-5s
Now go back to first equation and solve for s using the answer we got for t;
3s-4(8-5s)=14
3s-32+20s=14
23s=46
s=2
Now we have the answer for s, so plug that into equation 1 and find out what t=,
3(2)-4t=14
6-4t=14
-4t=8
t=-2
=)